A man walks westward on a ship with a speed of 4km/hr. The ship's propeller drives it 15km/hr south. What is the actual speed of the man relative to the Earth? What is the direction of his velocity?
1
Expert's answer
2021-10-28T08:52:13-0400

The actual speed is


"v=\sqrt{4^2+15^2}=15.52\text{ km/h}."

The direction is


"\theta=\tan^{-1}\frac{15}{4}=75°\text{ S of E}."